Fifty machines in a room produce a noise level of 100 dB (re 20 μPa). Using the exact method, answer the following: a. If 20 machines are removed, what will be the new noise level in the room? b. If in place of 20 machines, 25 machines are removed, calculate the new noise level. c. Would you recommend removing the additional five machines (part b)? Why or why not?
Answer #1

  50 machines produce 100 dB

so, (50-20)=30 machines produce 100*30/50=60 dB..........(a)

Again, additional 5 machines are removed then (60-10)=50 dB........(b)

(c) No, I did not recommend to removing the additional five machines.

Because, 50 dB is not harmful noise,It does not that much negative effect in nature and human